OpenStudy (anonymous):

How to parametrize the equation y^2 = x^3?

OpenStudy (cwrw238):

y = x^(3/2) x = y^(2/3) let x = t^2 then y = (t^2)^ 3/2 = t^3 and x = (t^3)^ 2/3 = t^2

OpenStudy (cwrw238):

so first find y in terms of x then x in terms of y and introduce an appropriate third parameter t write x in terms of t and substitute in the formula relating y and x (my last line wasn't really necessary as i had already defined x in terms of t but it served as a final check)

OpenStudy (zpupster):

i would like to add there is a nice discussion here http://www.physicsforums.com/showthread.php?t=14562

OpenStudy (anonymous):

Hmm, so let's say, we have \(x^{10} = y^7\) x = y^(7/10) y = x^(10/7) Let x = t^7, t= t^(10) In this case, aren't we just switching the power of x and y for the parameter t?